OpenStudy (love_333):

Medal + fan: A spherical balloon has a circumference of 19cm. What is the approximate surface area of the balloon to the nearest square cm? A 115cm^2 B 460cm^2 C 361cm^2 D 1,134cm^2

OpenStudy (love_333):

@Erak ?

OpenStudy (erak):

So first solve for your radius when given circumference. C = 2pi(r)

OpenStudy (love_333):

r=3.02 so 3

OpenStudy (love_333):

right?

OpenStudy (erak):

yeah, not rounding and leaving it as like 9.5/pi could help a little.

OpenStudy (erak):

So just stick that radius into the SA of a sphere formula SA = 4pi(r)^2

OpenStudy (love_333):

Okay with sticking to 9.5 I got 1,133.54

OpenStudy (erak):

oh it's "9.5/pi", so in your calc it's SA = 4pi(r)^2 SA = 4pi(9.5/pi)^2

OpenStudy (love_333):

If I do the first one i get: 1,133.54 If i do the second i get: 114.9098

OpenStudy (love_333):

Which is it..?

OpenStudy (erak):

the second one, r = 9.5/pi, which is equivalent to like 3.02 ish, just that having an exact radius will help. answer is A

OpenStudy (love_333):

Okay. Thank you!